How many atoms of O are present in 4.5 moles of MSO4? O A4 B. 1.08 x 1025 C. 18 D.2.71 x 1024
Answer #1

one mole of MgSO4 contains 4 moles of oxygen atom it means

1 mole of MgSO4 = 4 * 6.022 * 1023 oxygen atoms

4.5 mole of MgSO5 = 4.5 * 4* 6.022 * 1023 which is equals to = 1.08 * 1025 atoms

hence option B is correct